# ESC: Efficient Speech Coding with Cross-Scale Residual Vector Quantized Transformers

### v2

@@ 3.4 Cross-Scale Residual Vector Quantization
For efficient modeling of audio signals, ESC employs multi-scale features through a coarse-to-fine decoding process to generate bitstreams at different levels. Our codec integrates the more intuitive cross-scale residual vector quantization (CS-RVQ) structure proposed by Jiang et al. (2022a), eliminating the need for additional networks to fuse encoder and decoder features. As depicted in Algorithm 1, Algorithm 2 and Figure 1, the decoding process is conditioned on encoder features from various resolution scales. This distinguishes it from the commonly used residual vector quantization scheme that operates only at the bottom scale, which often overlooks low-level information.
Encoding. The encoding process begins with the encoder function $F_{\phi}(\cdot)$, creating multi-scale encoder features $\bm{z}_{e_{1}},...,\bm{z}_{e_{n}}$. $\bm{z}_{e_{n}}$ is first quantized by the bottleneck quantizer $Q_{0}$ to form the lowest bitstream. This represents the simplest case when $s$ is set to 1, and CS-RVQ reduces to a fixed-scale VQ at the bottleneck. At higher bitstreams, the residual between symmetric encoder and decoder features with higher resolutions, $\bm{z}_{e_{n-i+1}}-\bm{z}_{q_{i-1}}$, is quantized by $Q_{i}$. This quantized residual $\bm{q}_{i}$ is added back to $\bm{z}_{q_{i-1}}$ and decoded by the subsequent decoder layer $g_{\phi_{i}}(\cdot)$, producing the next decoder feature $\bm{z}_{q_{i}}$. Recursively, residuals at higher resolutions are progressively quantized, forming the remaining bitstreams. This mechanism enables multi-scale learning and allows the decoder layers to incrementally reduce quantization errors by conditioning on encoder features. When $s>2$, the encoding process of CS-RVQ requires $s-2$ additional decoder layers to produce residuals at higher levels. After encoding, $\bm{z}_{e_{0}}$ is compressed to be multi-level codes $\tilde{z}_{q_{0}},\tilde{z}_{q_{1}},\ldots,\tilde{z}_{q_{s-1}}$.
Decoding. The decoding process starts by retrieving the quantized latent from the bottom VQ using the code index $\tilde{z}_{q_{0}}$, which provides the original decoder input $\bm{z}_{q_{0}}$. At higher levels, $\tilde{z}_{q_{1}},...,\tilde{z}_{q_{s-1}}$ are iteratively used to retrieve codewords, resulting in multi-scale quantized residuals $\bm{q}_{1},...,\bm{q}_{s-1}$. Each quantized residual $\bm{q}_{i}$ is added back to the corresponding decoder feature $\bm{z}_{q_{i-1}}$ to improve the decoding process. Starting from the $s$-th decoder layer, there are no quantized residuals, and the layers perform vanilla decoding. Finally, the recovered frame spectrum $\bm{z}_{q_{n}}$ is obtained, which conditions on $s-1$ residual features.
Training. During the training of CS-RVQ codecs, the encoding and decoding processes are combined to form a complete forward pass. To enable bitrate scalability, we sample the number of bitstreams $s\sim\text{Uniform}\{1,\ldots,n\}$ at a rate $p$ within each training batch. $p$ is a hyperparameter that balances the reconstruction quality at different bitrates, as proposed by Kumar et al. (2023).
@@
1:Patch embedding frame $\bm{z}_{e_{0}}\in\mathbb{R}^{C_{0}\times H_{0}}$, encoder $F_{\phi}(\cdot)$, decoder $G_{\psi}(\cdot)$, vector quantizers $Q_{0},Q_{1},...,Q_{s-1}$, number of bitstreams $s$
@@
3:$\bm{z}_{q_{0}},\tilde{z}_{q_{0}}\leftarrow Q_{0}(\bm{z}_{e_{n}})$triangleright Bottleneck VQ
@@
7:end $\triangleright$ Encoding requires $s-2$ decoder layers
@@
2:$\bm{z}_{q_{0}}\xleftarrow[]{Q_{0}}\tilde{z}_{q_{0}}$triangleright Retrieve codewords from VQ
@@
6:end $\triangleright$ Decoding conditions on residuals
@@
9:end $\triangleright$ Continue with vanilla decoding
@@ 3.5 Mitigating Codebook Collapse
During quantization, the spectrum frame feature needs to be flattened, merging the frequency and channel dimensions. This results in large hidden dimensions for VQ, increasing the possibility of codebook collapse during training.
Vector quantization setups. To effectively optimize the codebooks, we modify vanilla VQs using a combination of product vector quantization and vector dimension reduction at each bitstream. Specifically, a frame feature $\bm{z}_{e_{i}}\in\mathbb{R}^{C_{i}\times H_{i}}$ is flattened into a $d$-dimensional vector before quantization. It is then split into a collection of $l$ sub-vectors. Each sub-vector $\bm{z}_{e_{i}}^{(m)}$ is down-projected and quantized by an individual codebook $\mathcal{C}_{m}$ to be ${\bm{z}}_{q_{i}}^{(m)}$:
$\bm{z}_{e_{i}}\equiv\{\bm{z}_{e_{i}}^{(m)}\mid\bm{z}_{e_{i}}^{(m)}\in\mathbb{R}^{d/l},m=1,...,l\},$ (8)
@@
Here $W_{\text{in}}\in\mathbb{R}^{\frac{d}{l}\times u}$, $W_{\text{out}}\in\mathbb{R}^{u\times\frac{d}{l}}$, and $u\ll d/l$. The projected vector $W_{\text{in}}^{\top}\bm{z}_{e_{i}}^{(m)}$ and codebook $\mathcal{C}_{m}$ are both $l_{2}$ normalized before finding the nearest codeword. This equalizes the scales of encoded vectors and codewords, enhancing codebook optimization by allowing a larger subset of codewords to receive gradient updates Łańcucki et al. (2020).
Pre-training paradigm. Training transformers can be challenging, and joint training with VQs is even harder. Therefore, we propose a pre-training paradigm that includes a warm-start to enhance the learning process. Initially, all VQ layers are deactivated, so no quantization occurs. Only the encoder and decoder are updated within the CS-RVQ framework, allowing latent features to bypass the quantizers and flow directly into the decoder layers. We term this the "pre-training" stage. Once the encoder and decoder have converged by minimizing reconstruction objectives, we proceed to train the entire VQ codec as usual. This approach helps mitigate the distribution shift of encoder features by optimizing the encoder in advance. It can stabilize codebook training and improve bitrate efficiency. Additionally, pre-training encoders and decoders is simpler, as it avoids the quantization errors associated with VQs. The detailed algorithm is illustrated in Appendix A.
@@ 3.6 Training Objectives
To train our codec, we employ a combination of reconstruction loss $\mathcal{L}_{recon}$ and vector quantization loss $\mathcal{L}_{vq}$. The reconstruction loss, $\mathcal{L}_{recon}$, consists of two components: an $l_{2}$ distance between the complex spectrum $\mathcal{X}$ and its reconstruction $\hat{\mathcal{X}}$, weighted by $\lambda_{1}$, and a multi-scale mel-spectrogram loss (Kumar et al., 2023), weighted by $\lambda_{2}$. These are denoted as $\mathcal{L}_{stft}$ and $\mathcal{L}_{mel}$, respectively:
@@
The vector quantization loss, $\mathcal{L}_{vq}$, comprises standard codebook and commitment losses as described in Eq. 2. It is averaged within the $l$ product VQs at each bitstream and summed across $s$ transmitted bitstreams. The final objective for end-to-end training is the summation of $\mathcal{L}_{recon}$ and $\mathcal{L}_{vq}$.
Figure 2: Performance evaluation results of different baseline codecs: dashed lines represent DAC baselines and solid lines represent our ESC models, with x-axis being transmission bits per second and y-axis being PESQ ()($\uparrow$)( ↑ ), Mel-Distance ()($\downarrow$)( ↓ ) and SI-SDR()($\uparrow$)( ↑ ). The metrics are averaged over our composed 1158 10-second speech clips.
@@ 4.1 Experimental Setup
Datasets. We extract 150 hours 16kHz multilingual clean speech from the DNS Challenge 2021 dataset Reddy et al. (2021). Training samples are clipped into 3-second segments, and validation samples are clipped into 10-second segments. For evaluation, we compile 1158 multilingual 10-second speech clips from the LibriSpeech Panayotov et al. (2015), Multilingual LibriSpeech Pratap et al. (2020), and AIShell Shi et al. (2020) datasets. The speakers are non-overlapping in each split.
Baselines. We compare our codec against the current state-of-the-art time-domain codec DAC, by reproducing three versions [^2] Reproduction settings are detailed in Appendix B.1. on our dataset:
1) DAC-Base (Adversarial): Descript’s original released codec, operating on 16kHz audio signals. It has 74M parameter count in total.
@@
3) DAC-Tiny (Non-Adversarial): A smaller and non-adversarial version of DAC to assess the impact of discriminator on improving audio fidelity.
@@
1) ESC-Base (Non-Adversarial): This base version consists of 6 encoder/decoder layers, with bitrates ranging from 1.5 to 9.0 kbps, and contains 8.39M parameters when operating at 9.0 kbps.
2) ESC-Base (Adversarial): Trained adversarially using the same multi-scale waveform and spectrogram discriminator as DAC.
3) ESC-Large (Non-Adversarial): A scaled-up version of ESC-Base with increased Swin Transformer layer depth, resulting in 15.58M parameters when operating at 9.0 kbps.
The ESC variants are trained using the AdamW optimizer with a learning rate of 1e-4 and betas set to (0.9, 0.999). Training runs up to 0.4 million iterations without learning rate schedulers. The proposed pre-training phase consists of 0.75 million iterations. After pre-training, the codebooks are initialized with a Kaiming normalization distribution He et al. (2015). The quantization dropout rate is set to 0.75. Loss weighting hyperparameters are set as $\lambda_{1}=0.25$, $\lambda_{2}=1.0$, and the commitment loss weighting $\beta=0.25$. In ESC-Base (Adversarial), the $\mathcal{L}_{stft}$ component is replaced with the $l_{1}$ feature matching loss, following DAC’s approach. The adjusted loss weights are $\lambda_{1}=15.0$, $\lambda_{2}=2.0$, and $\beta=0.25$.
Automatic evaluation metrics. We use objective metrics to efficiently evaluate reconstruction performance. These include the PESQ score Union (2007) from the speech enhancement domain, following Jiang et al. (2022a), the $L_{1}$ distance between log mel-spectrograms of reference and decoded waveforms (mel distance) Kumar et al. (2023), and the scale-invariant source-to-distortion ratio (SI-SDR) Le Roux et al. (2019). To measure codec inference latency, we use the real-time factor (RTF), defined by Défossez et al. (2023) as the ratio of speech audio duration to model processing time.
4.2 Comparison with DAC
We present a comprehensive comparison focusing on compression performance and inference efficiency, as demonstrated in Figure 2 and Table 1.
Performance evaluation. First, note ESC-Base and DAC-Tiny are similar in model size, each with approximately 8 million trainable parameters. Our results show that ESC-Base consistently outperforms DAC-Tiny across all bitrates, even without an adversarial discriminator. In contrast, DAC-Tiny’s reconstruction quality drops significantly without a discriminator, particularly in SI-SDR statistics. This indicates a heavy reliance of DAC models on GANs for maintaining high reconstruction quality. Notably, ESC-Base is compatible with the same convolution-based GAN discriminator used in DAC, as evidenced by its improved performance across all metric curves in its adversarial variant. Besides, ESC-Large demonstrates that increasing ESC’s model size can further enhance performance, with its PESQ curve matching that of DAC-Base, the top-performing and largest model. While DAC-Base achieves higher SI-SDR values, ESC-Large has a smaller mel-distance. Thus, we conclude that the two codecs achieve comparable performance, even though ESC-Large is trained without GANs.
@@ 4.2 Comparison with DAC
Complexity evaluation. Despite their exceptional performance, Descript’s top-performing codecs face significant computational challenges. This is evident from Table 1, where the decoding real-time factor (RTF) for DAC-Base is approximately 3.0, making it impractical for real-time applications. In contrast, our transformer-based ESC achieves much higher RTFs, indicating superior computational efficiency. Although ESC-Base is not as fast as DAC-Tiny due to the overhead of attention computation, it offers substantially better speech reconstruction capabilities, striking a favorable balance between compression performance and latency. Moreover, following the CS-RVQ scheme, ESC possesses enhanced encoding speeds at lower bitrates—a capability not found in DAC models.
@@
These results suggest that our transformer-based codec, equipped with CS-RVQ, is a more parameter-efficient foundational codec compared to time-domain convolutional codecs. We can conclude that ESC is a more lightweight and effective neural codec, because ESC-Large achieves comparable performance to DAC-Base without requiring a powerful GAN discriminator, with approximately $\times$4.8 smaller model size, $\times$1.4 times encoding speed, and $\times$6.4 faster decoding speed.
4.3 Ablation Study
To investigate the effectiveness of the proposed components in ESC, we conducted thorough ablation experiments [^4] Implementation setups are detailed in Appendix B.3. by training frequency-domain codecs with different architectures. For fair comparisons, all other ablation models listed in Table 2 have similar model sizes to ESC-Base.
Swin Transformers and CNNs. To demonstrate that transformers are superior backbones for encoder/decoder in speech coding, we focus on two pairs of experiments: CNN/SwinT + RVQ and CNN/SwinT + CS-RVQ. In these experiments, the channel dimensions of the CNN blocks are set to match the hidden dimensions of the Swin Transformer Blocks (STBs). The comparison, as shown in the results, reveals that transformer-based codecs consistently outperform CNN-based codecs across all performance metrics and bitrates, regardless of the quantization scheme used.
CS-RVQ and RVQ. Table 2 also highlights that CS-RVQ is a superior quantization scheme compared to RVQ, regardless of whether the backbone is CNN or STB. RVQ-based codecs would hit performance bottlenecks, as adding more VQs does not improve audio quality (e.g., from 6.0kbps to 9.0kbps). However, codecs using the CS-RVQ scheme do not face such bottlenecks at higher bitrates and consistently outperform their RVQ counterparts. CS-RVQ is thus a more efficient scheme that leverages multi-scale features effectively.
Effect of pre-training paradigm. To evaluate the efficacy of a pre-training stage, we conducted an experiment of ESC-Base w/o pre-training. We monitored the VQ utilization rate, calculated as the sum of entropy (in bits) divided by the maximum number of bits from all transmitted bitstreams. This metric reflects bitrate efficiency and the fraction of seldom-used codewords. The results indicate that all experiments with pre-training achieve a near 1.0 utilization rate. However, ESC-Base w/o pre-training displays a lower utilization rate at 9.0kbps, and its reconstruction performance is also worse than ESC-Base. These findings suggest that including a pre-training stage indeed helps avoid bitrate wastage and improve audio reconstruction quality.

### v3

@@ 3.4 Cross-Scale Residual Vector Quantization
To achieve parameter-efficient modeling of audio signals, ESC employs multi-scale features that capture coarse-to-fine information. It integrates the more intuitive residual-based cross-scale vector quantization (CS-RVQ) framework proposed by Jiang et al. (2022a), eliminating the need for additional networks to merge encoder and decoder features for improved reconstruction quality. As depicted in Algorithm 1, Algorithm 2 and Figure 1, the decoding process is conditioned on the encoded quantized residuals between encoder and decoder features from low-to-high resolution scales. This approach differs from the commonly used residual vector quantization scheme, which operates solely at the lowest scale, relying on the highest-level information while overlooking low-level details.
Encoding. The encoding process begins with the encoder $F_{\phi}(\cdot)$, creating multi-scale encoder features $\bm{z}_{e_{1}},...,\bm{z}_{e_{n}}$. $\bm{z}_{e_{n}}$ is first quantized by the bottleneck quantizer $Q_{0}$ to form the lowest bitstream. This represents the simplest case when the number of transmitted bitstream $s$ is set to 1, and CS-RVQ reduces to a fixed-scale VQ at the bottleneck. For higher bitstreams, the residual between symmetric encoder and decoder at higher resolutions, $\bm{z}_{e_{n-i+1}}-\bm{z}_{q_{i-1}}$, is quantized by $Q_{i}$. The quantized residual $\bm{q}_{i}$ is then added back to $\bm{z}_{q_{i-1}}$ and decoded by the subsequent decoder layer $g_{\phi_{i}}(\cdot)$, producing the next decoder feature $\bm{z}_{q_{i}}$. Recursively, residuals at higher resolutions are progressively quantized, forming the remaining bitstreams (see Algorithm 1, Lines 3-6). This mechanism enables multi-scale learning, allowing the decoder layers to incrementally reduce quantization errors by conditioning on encoder-decoder residual features. When $s>2$, this encoding process requires forward passing $s-2$ additional decoder layers to produce residuals at higher levels. After encoding, the input $\bm{z}_{e_{0}}$ is compressed into multi-level codes $\tilde{z}_{q_{0}},\tilde{z}_{q_{1}},\ldots,\tilde{z}_{q_{s-1}}$.
@@
1:A flattened time frame $\bm{z}_{e_{0}}\in\mathbb{R}^{C_{0}H_{0}}$, encoder $F_{\phi}(\cdot)$, decoder $G_{\psi}(\cdot)$, vector quantizers $Q_{0},Q_{1},...,Q_{s-1}$, number of bitstreams $s$
@@
3:$\bm{z}_{q_{0}},\tilde{z}_{q_{0}}\leftarrow Q_{0}(\bm{z}_{e_{n}})$triangleright bottom VQ
@@
7:end $\triangleright$ Encoding involves $s-2$ decoder layers
@@
2:$\bm{z}_{q_{0}}\xleftarrow[]{Q_{0}}\tilde{z}_{q_{0}}$triangleright Retrieve codewords from bottom VQ
@@
6:end $\triangleright$ Decoding refined by quantized residuals
@@
9:end $\triangleright$ Continue with regular decoding
@@
Decoding. The decoding process starts by retrieving the quantized latent at the bottom VQ using code $\tilde{z}_{q_{0}}$, which provides the initial decoder input $\bm{z}_{q_{0}}$. At higher levels, the codes $\tilde{z}_{q_{1}},...,\tilde{z}_{q_{s-1}}$ are iteratively used to retrieve codewords, producing multi-scale low-to-high quantized residuals $\bm{q}_{1},...,\bm{q}_{s-1}$. In Algorithm 2, Lines 2-5, each quantized residual $\bm{q}_{i}$ is added back to the corresponding decoder feature $\bm{z}_{q_{i-1}}$ to refine the decoding process. Starting from the $s$-th decoder layer, there are no quantized residuals, and the remaining layers perform regular decoding. Finally, the recovered frame vector $\bm{z}_{q_{n}}$ is obtained, benefiting from $s-1$ quantized residual features.
Training. During training, the encoding and decoding processes are concatenated to form a complete forward pass. To enable bitrate scalability, we sample $s\sim\text{Uniform}\{1,\ldots,n\}$ at a rate $p$ within each training mini-batch. $p$ is a hyperparameter that balances the reconstruction quality at different bitrates, as proposed by Kumar et al. (2023).
@@ 3.5 Mitigating Codebook Collapse
ESC performs a per-frame vector quantization. Before nearest neighbor searching, each input spectrum frame feature in $\mathcal{Z}$ needs to be flattened, merging the frequency and channel dimensions. This approach can result in large input vector dimensions for VQ, increasing the optimization challenges associated with codebook underutilization.
Vector quantization setups. To optimize the codebooks effectively, we modify the vanilla VQ by combining product vector quantization with code-vector factorization at each bitstream. Specifically, a flattened $d$-dimensional frame vector $\bm{z}_{e_{i}}$ is split into a set of $l$ sub-vectors. Each sub-vector $\bm{z}_{e_{i}}^{(m)}$ is down-projected by $W_{\text{in}}\in\mathbb{R}^{\frac{d}{l}\times u}$, where $u\ll d$, and then quantized using an individual codebook $\mathcal{C}_{m}$. The selected code-vector is then up-projected by $W_{\text{out}}\in\mathbb{R}^{u\times\frac{d}{l}}$ to form ${\bm{z}}_{q_{i}}^{(m)}$:
$\bm{z}_{e_{i}}\equiv\{\bm{z}_{e_{i}}^{(m)}\mid\bm{z}_{e_{i}}^{(m)}\in\mathbb{R}^{\frac{C_{i}H_{i}}{l}},m=1,...,l\},$ (8)
@@
Additionally, both the projected vector $W_{\text{in}}^{\top}\bm{z}_{e_{i}}^{(m)}$ and codebook $\mathcal{C}_{m}$ are $l_{2}$ normalized before computing the distance matrix. This equalizes the scales of input vectors and codewords, enhancing codebook optimization by allowing a larger subset of codewords to receive gradients Łańcucki et al. (2020).
Pre-training paradigm. Training transformers can be challenging, and jointly training them with VQ layers is even more difficult. To address this, we propose a pre-training paradigm that includes a warm-start to facilitate the learning process. Initially, all VQ layers are deactivated, meaning no quantization occurs. During this "pre-training" stage, only the encoder and decoder are updated within the CS-RVQ framework, allowing latent features to bypass the quantizers and flow directly into the decoder layers. Once the encoder and decoder have converged by minimizing reconstruction objectives, we resume training the entire VQ codec as usual. This approach helps mitigate the distribution shift of encoder representations by pre-optimizing the encoder. It helps stabilize codebook training and improve bitrate efficiency. Moreover, pre-training an auto-encoder is simpler, as it avoids the quantization errors associated with VQs. The detailed algorithm is provided in Appendix A.
@@ 3.6 Training Objectives
To train our codec, we use a combination of reconstruction loss $\mathcal{L}_{recon}$ and vector quantization loss $\mathcal{L}_{vq}$. The reconstruction loss, $\mathcal{L}_{recon}$, consists of two components: an $l_{2}$ distance between the complex spectrum $\mathcal{X}$ and its reconstruction $\hat{\mathcal{X}}$, which forces the model to reconstruct the real and imaginary parts, weighted by $\lambda_{1}$, and a multi-scale mel-spectrogram loss (Kumar et al., 2023), weighted by $\lambda_{2}$. These are denoted as $\mathcal{L}_{stft}$ and $\mathcal{L}_{mel}$:
@@
$\mathcal{L}_{vq}$ comprises the standard codebook and commitment losses as described in Equation 2. It is averaged across the $l$ product vector quantizers and summed over all $s$ bitstreams. The final objective for joint optimization is the summation of $\mathcal{L}_{recon}$ and $\mathcal{L}_{vq}$. To deactivate the VQ layers during the pre-training stage, $\mathcal{L}_{vq}$ is set to zero.
Figure 2: Reconstruction quality evaluation of different baseline codecs: dashed lines represent DAC baselines and solid lines represent our ESC models, with x-axis being transmission bits per second and y-axis being PESQ ()($\uparrow$)( ↑ ), Mel-Distance ()($\downarrow$)( ↓ ) and SI-SDR()($\uparrow$)( ↑ ). The metrics are averaged over our composed 1158 10-second speech clips.
@@ 4.1 Experimental Setup
Datasets. We extract 150 hours of 16kHz multilingual clean speech from the DNS Challenge dataset Reddy et al. (2021). Training samples are clipped into 3-second segments, and validation samples into 10-second segments. For evaluation, we compile 1158 multilingual 10-second speech clips with non-overlapping speakers from the LibriSpeech Panayotov et al. (2015), Multilingual LibriSpeech Pratap et al. (2020), and AIShell Shi et al. (2020) datasets.
Baselines. We compare our ESC against the current state-of-the-art time-domain codec DAC, by reproducing three versions [^2] Reproduction settings are detailed in Appendix B.1. on our dataset:
1) DAC-Base (adversarial): Descript’s original released codec, operating on 16kHz audio signals. It has 74M parameter count in total. Its associated discriminator has 42M additional parameter count.
@@
3) DAC-Tiny (non-adversarial): A smaller and non-adversarial version of DAC to assess the impact of discriminators on improving audio fidelity.
@@
1) ESC-Base (non-adversarial): A base version codec consisting of 6 encoder/decoder layers, with bitrates ranging from 1.5 to 9.0 kbps. It contains 8.39M parameters when operating at 9.0 kbps.
2) ESC-Base (adversarial): An adversarial version using the same multi-scale multi-band waveform and spectrogram discriminator in DAC.
3) ESC-Large (non-adversarial): A scaled-up version with increased Swin Transformer layer depth, having 15.58M parameters at 9.0 kbps.
Our ESC variants are trained using the AdamW optimizer Loshchilov (2017) with a learning rate of 1e-4 and a weight decay of 1e-2. Training runs up to 0.4 million iterations without learning rate schedulers. The proposed pre-training phase consists of 0.75 million iterations. After pre-training, the codebooks are initialized with a Kaiming normalization distribution He et al. (2015). The quantization dropout rate $p$ is set to 0.75. Loss weighting hyperparameters are set as $\lambda_{1}=0.25$, $\lambda_{2}=1.0$, and the commitment loss weighting $\beta=0.25$. For ESC-Base (adversarial), the $\mathcal{L}_{stft}$ component is eliminated. We use the HingeGAN Lim and Ye (2017) adversarial loss formulation and the $l_{1}$ feature matching loss Kumar et al. (2019), following the approach of DAC.
Automatic evaluation metrics. We use objective metrics to efficiently evaluate reconstruction performance. These include the PESQ score Union (2007) from the speech enhancement domain, following Jiang et al. (2022a); the $l_{1}$ distance between log mel-spectrograms of reference and decoded waveforms (Mel-Distance) Kumar et al. (2023); and the scale-invariant source-to-distortion ratio (SI-SDR) Le Roux et al. (2019). To measure codec inference latency, we use the real-time factor (RTF), defined as the ratio of speech audio duration to model processing time Défossez et al. (2023).
@@
4.2 Comparison with DAC
We provide a thorough comparison focusing on compression rate, reconstruction quality, and inference efficiency, as shown in Table 1 and Figure 2.
Performance evaluation. First, it is important to note that ESC-Base and DAC-Tiny are similar in model size, each with approximately 8 million trainable parameters. Our results show that ESC-Base consistently outperforms DAC-Tiny across all bitrates, even without an adversarial discriminator. In contrast, DAC-Tiny’s reconstruction quality significantly drops without a discriminator in training, particularly in SI-SDR statistics. This indicates a heavy reliance of DAC models on GANs for maintaining high reconstruction quality. Notably, ESC-Base is compatible with the same convolution-based GAN discriminator used in DAC, as evidenced by its improved performance across all metric curves in its adversarial variant. Additionally, ESC-Large demonstrates that increasing ESC’s model size can further enhance performance, with its PESQ curve matching that of DAC-Base, the top-performing and largest model. While DAC-Base achieves higher SI-SDR values, ESC-Large records a smaller Mel-Distance. Thus, we conclude that the two codecs achieve comparable performance, even though ESC-Large is trained without an adversarial discriminator.
@@ 4.2 Comparison with DAC
Complexity evaluation. Despite their exceptional performance, Descript’s top-performing codecs face significant computational challenges. This is evident from Table 1, where the decoding real-time factor (RTF) for DAC-Base is approximately 3.0, making it rather impractical for real-time applications. In contrast, our transformer-based ESC achieves much higher decoding RTFs (approximately 34), indicating superior computational efficiency. Although ESC-Base is not as fast as DAC-Tiny due to the overhead of attention computation, it offers substantially better speech reconstruction capabilities, striking a favorable balance between compression performance and computation latency. Future work could incorporate transformer speedup techniques such as FlashAttention Dao et al. (2022) to further enhance ESC’s latency further. Moreover, following the CS-RVQ scheme, ESC possesses faster encoding speeds at lower bitrates—a capability not evidently found in DAC models.
@@
These results suggest that our transformer-based codec, equipped with CS-RVQ, is a more parameter-efficient foundation model compared to time-domain convolutional counterparts. ESC is shown to be a more lightweight and effective neural speech codec, as ESC-Large achieves comparable performance to DAC-Base without the need for a powerful discriminator. Specifically, it boasts approximately $\times$4.8 smaller model size, $\times$1.4 faster encoding speed, and $\times$6.4 faster decoding speed.
@@
4.3 Ablation Study
To investigate the effectiveness of the proposed components in ESC, we conducted thorough ablation experiments [^4] Implementation setups are detailed in Appendix B.3. by training frequency-domain codecs operating on complex STFT spectra with different architectures. For fair comparisons, all other ablation models listed in Table 2 have similar model sizes to ESC-Base.
Swin Transformers and CNNs. To demonstrate that transformers are superior auto-encoder backbones in neural speech coding, we focus on two pairs of experiments: CNN/SwinT + RVQ and CNN/SwinT + CS-RVQ. In these experiments, the channel dimensions of the CNN blocks are set to match the hidden dimensions of the Swin Transformer Blocks (STBs). The comparison, as shown in Table 2, reveals that transformer-based codecs consistently outperform CNN-based codecs across all performance metrics and bitrates, regardless of the quantization scheme used.
CS-RVQ and RVQ. Table 2 highlights that CS-RVQ is a superior quantization scheme compared to RVQ, regardless of whether the backbone is CNN or STB. RVQ-based codecs hit performance bottlenecks, as adding more VQs does not improve audio quality (e.g., from 6.0 kbps to 9.0 kbps). However, codecs using the CS-RVQ scheme do not face such bottlenecks at higher bitrates and consistently outperform their RVQ counterparts. CS-RVQ is therefore a superior vector quantization framework that leverages multi-scale features effectively.
Effect of pre-training paradigm. To evaluate the efficacy of the pre-training stage, we conducted an experiment of ESC-Base w/o pre-training. We monitored the VQ utilization rate, calculated as the sum of entropy (in bits) divided by the maximum number of bits from all transmitted bitstreams. This metric reflects bitrate efficiency and the fraction of seldom-used codewords. The results indicate that models with pre-training achieve a near 1.0 utilization rate. However, ESC-Base w/o pre-training displays a lower utilization rate at 9.0 kbps, and its reconstruction performance is also inferior to that of the fully pre-trained ESC-Base. These findings suggest that the pre-training paradigm indeed helps avoid bitrate wastage and improve audio reconstruction quality.